# On The Computational Complexity for Minimizing Aerial Photographs for Full Coverage of a Planar Region

### v1

@@ Abstract
With the popularity of drone technologies, aerial photography have become prevalent in many daily scenarios such as environment monitoring, structure inspection, law enforcement etc. A central challenge in this domain is the efficient coverage of a target area with photographs that can entirely capture the region, while respecting constraints such as the image resolution, and limited number of pictures that can be taken. This work investigates the computational complexity of several fundamental problems arised from this challenge. By abstracting the aerial photography problem into the coverage problems in computational geometry, we demonstrate that most of these problems are in fact computationally intractable, with the implication that traditional algorithms cannot solve them efficiently. The intuitions of this work can extend beyond aerial photography to broader applications such as pesticide spraying, and strategic sensor placement.
@@ Introduction
Consider the scenario when a drone is tasked to take pictures to capture the boundary of some cropland with a gimbal camera attached to it. We can change the zoom factor $\lambda$ of the camera to make the image footprint larger or smaller with the same resolution, e.g. 3000 x 4000 in 4K resolution. The gimbal camera usually faces downwards when the drone flies high in the sky to take pictures for terrain inspection for applications like agricultural analysis or mapping. With a greater zoom factor of the gimbal camera, the projected footprint on the ground becomes smaller, but the picture taken contains more detail for each part on the ground.
Unfortunately, the computing or disk resource on the drone is usually limited, so let’s assume the drone can only take at most $k$ pictures in one flight, where $k$ is a parameter induced by the drone hardware. We want to ask the question that what is the largest zoom factor $\lambda$ we can give the gimbal camera such that we can still capture all parts of the region with $k$ pictures. Figure  1 shows a coverage result for covering the boundary or interior of a region with circles or squares, where cirlces can be imagined as footprints of visible regions using downwards facing fisheye cameras or radars, and squares as the specialization of rectangular images taken by regular cameras.
@@
With the ease of obtainability of drone hardware, especially quadrotors, vast aerial applications were developed, for environmental monitoring of static areas (manfreda2018use), 3D scanning (Roberts2017Submodular), and more complicated multi aerial/ground vehicle coordination case (grocholsky2006cooperative; pitla2020ground; pretto2020building).
The theoretical aspect of the aerial photo taking is strongly related to area of computational geometry, where coverage of point sets (aka. geometric pointclouds) are analyzed to a great extent. For example, the $p$-center problem asks that given a set of points, and a number $p$, what is the minimum radius of $p$ circles to cover all the points (toth2017handbook). Similar problems include the $k$-median problem that asks to minimize the summation of the distances to the closest center points, and the $k$-means problem that tries to solve the sum of squared distances from each point to the closest point among the $k$ chosen points. These problems, especially $k$-means, can go to high dimension in machine learning. However, for the real-world coverage problems for aerial phography, the point sets are associated with specific physical meanings, and usually lie in the $3$-dimensional Euclidean space in the simple case, or in $5$ dimensions when normal directions are included.
Notably, the majority of the interesting regions to cover in aerial coverage problems have some extent of continuity, for example, a complete structure, a continuous boundary or fence. Geometric shapes like 2d polygons or 3d meshes create more challenges for the usually studied point set coverage problems while making it more realistic. In robotics, this characteristic is usually implied, as a robot handling the coverage or inspection task moves in a continuous manner, and it makes sense to assume the region is continuous instead of being a scattered point set.
Early studies on covering and packing (Hachbaum1985Approximation) provide strong hardness results and approximation scheme for covering points. Later work (feder1988Optimal) further showed $1.88$-inapproximability gap for point sets coverage using $L_{2}$ metric, and $2$-inapproximability gap for coverage using $L_{1}$ metric. For continuous regions and boundarys of a simple polygon, recent work (Feng2020Optimally) has shown $1.152$-inapproximability gap with the minimum circle radius using $k$ circles.
@@
In this paper, the complexity analysis for covering a simple planar polygon using squares is the main focus. Specifically, this work shows the inapproximability gap of $1.165$ for finding the minimum side length of $k$ squares to cover the simple 2D region (a simple polygon) and the inapproximability gap of $1.25$ when the square centers are restricted to be inside the region. This work also develops a $2.828$-optimal algorithm to solve the square coverage problem in polynomial time. Since most regular cameras used for aerial photography take rectangular images, the results in this work bring the computational complexity results shown in (Feng2020Optimally) closer to reality. And it can potentially be a reference for setting cameras’ zooming factors for aerial imaging with regular cameras. Additionally, as the robot’s position may be restricted to be only on top of the region to take pictures for the region itself, the paper also considers the variant of the problem when the positions of the square or circle centers must be inside the simply connected region or on top the region’s perimeter. This give the work in (Feng2019Efficient) an aerial photography interpretation.
@@
Given a simple polygon (a planar polygonal chain without self-intersections or holes) $P$, and a number $k\geq 1$. What is the minimum length $\ell$ such that we can put $k$ circles each with radius of $\ell$, such that $P$ can be contained by these circles.
@@
Given a simple polygon (a planar polygonal chain without self-intersections or holes) $P$, and a number $k\geq 1$. What is the minimum length $\ell$ such that we can put $k$ squares each with side length of $\ell$, such that $P$ can be contained by these squares.
@@
If we only care about covering the boundary of a simple polygon, similar to the definition of the barrier coverage problem (Gage1992Command). This problem can be reduced to the two problems aforementioned because we can solidify the boundary by expanding the edges to sticks with a small width $\delta$, and leave a narrow opening on the polygon. When $\delta\rightarrow 0$, the results for covering the solidified boundary skeletion is the same as the results for covering just the boundary skeletion. In this way, covering the expanded polygon from the original perimeter is the same as covering the boundary of the original polygon. Conversely, a simple polygon can be approximately represented by a polygon whose boundary almost fills in the interior of the polygon. Hence, the computational complexity of covering the interior and the boundary of a simple region is at the same level.
In many real scenarios, the sensors or the robots (i.e. the centers of the circles or squares) must be inside the region itself or on the perimeter, for example, the watch towers should be built on the defensive wall itself and not outside; unmanned aerial vehicles for securities should not go outside or inside the security zone lines. When the constraints that the circle or square centers must lie inside the polygon must be satisfied, the constrained version of Problem 2 and  1 are called.
Problem 3 (Constrained Square or Circle Coverage)
Same as Problem  1 and Problem  2, but the circle or square centers must lie inside the polygon $P$.
@@
Similarly, the case when the feasible region to place circle centers is the interior of the polygon contains the case when the feasible region is on the boundary of a polygon, because we can easily make a polygon out of the boundary of a simple polygon by expanding the line segments to some certain width, and make narrow openings for holes elimination.
@@
Given a planar cubic graph $G(V,E)$, we convert it into a special structure through a sequence of steps. The first step embeds the graph $G$ into the plane. Each vertex in $G$ becomes a vertex junction, and each edge becomes a “fence” like structure connecting two neighboring junctions which is an odd length path intersected with a set of perpendicular bars with length $\zeta$ ($\zeta$ is a parameter to be determined for different proofs), at each unit line segment.
@@
From here, a second modification is made which completes the construction of $T_{G}$. In each previously constructed path $u\cdots w=uv_{1}\ldots v_{2m}w$, for each $v_{i}v_{i+1}$, $1\leq i\leq 2m-1$, we add a line segment of length $\zeta$ that is perpendicular to $v_{i}v_{i+1}$ such that $v_{i}v_{i+1}$ and the line segment divide each other in the middle. $G^{\prime}$ and the bars form $T_{G}$.
@@
The theorem in this section is proved completely in (Feng2020Optimally), but for the completeness of this paper and the arguments in the later parts, we provide a sketch of the proof.
@@
The gadget structure (refer it as $T$) constructed in the previous section can be easily converted to a simple polygon by dilating the lines to a thin structure with a small width $\varepsilon$ while leaving a narrow opening with width $\varepsilon$ to make it a simple polygon (refer it as $P_{\epsilon}$). Assume the minimum radius of $k$ circles to cover a simple polygon is $\ell^{*}$, then the minimum radius $\ell$ for covering $P_{\epsilon}$ must satisfy $\ell-\varepsilon\leq\ell^{*}\leq\ell+\epsilon$. When $\epsilon\rightarrow 0$, $\ell$ and $\ell^{*}$ will converge to be the same. So, we only need to show the complexity of approximation to put $k$ circles with min radius to cover the structure $T$.
@@
Since each junction maps to a vertex in $G$ in the transformation, call the side of the edge with a circle covering only one bar the “odd-end”, and the other side of the edge the “even-end”. If there is a vertex cover of graph $G$, then it is possible to let all the vertex been chosen be the “even-end” or “odd-end”, and the rest of the vertices be “even-end” in the edge link coverage pattern. Since all the junction represented vertex either inside a vertex cover or has a neighbor inside the vertex cover, it is always possible to find such a “odd-end” and “even-end” combination for the junctions so at least no junction has all three edges been even ends.
On the other side, for a coverage of the special structure with the minimum number of circles, select those vertices with at least one even-end. It is clear that these vertices form a vertex cover.
@@
Listing all the geometric equations in Figure  4 gives,
@@
Integrating the system of equations from (1) to (6) together gives $\ell=1.152$, which is the extreme case when the circle coverage pattern of $T_{G}$ cannot map to a vertex cover of the orginal vertex cover problem.
@@
In the structure gadget constructed in the previous section, we shrink the unit length segment to $\sqrt{2}/2$, so the current distance between neighboring bars is set to be ${\sqrt{2}}/{2}$. And the bar length $\zeta$ here is set to be $\zeta={\sqrt{2}}/{2}$.
@@
Figure 5: Structure within the odd length path, where each unit length segment is now shrinked to $\sqrt{2}/2$, and the vertical bar length is $\zeta=\sqrt{2}/2$.
@@
Figure 6: Possible square coverage pattern along the constructed odd-length path until it reaches $1.25$.
@@
Then, let’s prove when there is $n+M$ squares with edge length less than $1.165$ to cover the tri-net structure, there is a vertex cover with of size $n$ for the graph $G$.
@@
Figure 8: The pattern of covering the crossing in the special structure constructed at the extreme square side length.
@@
At the maximum square side length of $\ell$ such that the structure still needs the same number of squares to cover as using unit squares, Figure  8 shows this situation when the junction can be covered with three squares among which two squares each covers one vertical bar and one square covers two vertical bars.
Listing all the geometric equations in Figure  8 gives,
@@
This inapproximability factor can be interpreted as the maximum level of zoom factor the camera mounted on the drone can have before taking pictures, or the minimum Ground Sampling Distance (GSD) the $k$ pictures can provide subject to full coverage of the boundary of a simple region using a polynomial time algorithm unless P=NP.
@@
In the construction and also in the problem setting in Problem  1 and  2, it is assumed that the center of the circle or square can be anywhere on the plane. However, in many robotics scenarios, the center of the circles or squares, i.e. the robot locations, have restrictions like they must be on top of the perimeter or inside the region. If those restrictions are applied, are these problems still NP-hard or they can be proved to have a higher inapproximability ratio?
Since if there is a solution of $k$ unit circle to cover the simple polygon constructed, they can be made to be centered at the skeletion structure $T_{G}$ constructed before until the circle radius of $\ell$ reaches $1.152$, the restricted version of the problem have at least the same inapproximability ratio.
@@
At the maximum square side length of $\ell$ such that the structure still needs the same number of squares to cover as using unit squares, Figure  9 shows this situation when the junction can be covered with three squares among which two squares covering one vertical bar each and one square covering two vertical bars.
Listing all the geometric equations in Figure  9 gives,
@@
Integrating the system of equations in (14) to (17) together gives $\ell=1.289$. Considering the possible change of coveage pattern along the odd-length-path depicted in Figure  6 when $\ell$ reaches $1.25$. We have,
@@
Compared to unrestricted version of the square coverage problem which has the inapproximability ratio of $1.165$, the inapproximability ratio for the restricted version for the square coverage problem is higher by around $8.5\%$.
@@
In a previous work (Feng2020Optimally), sampling followed by applying the traditional $2$-optimal farthest clustering algorithm (Gonzalez1985Clustering) for the $k$-center problem gives a $(2+\epsilon)$-optimal result. When replacing the farthest clustering algorithm with mathematical programming, with the state of the art integer programming tool (Optimization2019Gurobi), a $(1+\epsilon)$-optimal result for $k\leq 100$ can be obtained in $1$ minute on a desktop platform with intel i7 and 16GB memory.
For the square coverage problem, which is similar to the $k$-center problem with $L_{\infty}$ distance metric where the distance between two points $p_{1}=(x_{1},y_{1})$ and $p_{2}=(x_{2},y_{2})$ is $\max(\|x_{1}-x_{2}\|,\|y_{1}-y_{2}\|)$. If we are using $L_{\infty}$ distance metric, which means the square used for covering the target region is always axis-aligned, using the same farthest clustering algorithm (Gonzalez1985Clustering) can produce a 2-optimal solution. Besides, a square with side length $\ell$ facing at any direction can be fully covered by an axis-aligned square with side length $\sqrt{2}\ell$. So, squares created with regular farthest clustering algorithm for the sampled points using the $L_{\infty}$ metric gives a $(2\sqrt{2}+\epsilon)\approx(2.828+\epsilon)$-approximation algorithm.
In the constrained version of the square and circle coverage problems, since the farthest clustering algorithm only select points (circle or square centers) among the sampled points in the target region or boundary, the results of the previous algorithm automatically satisfy the constraints. Thus, the efficient algorithms for the square and circle coverage problems still apply for approximately solving the constriained circle and square coverage problem within an approximation factor of $2$ and $2.828$, respectively.
@@
1 Sample NN points from the polygon
2 Run the 22-optimal farthest clustering algorithm with $L_{1}$ metric to get center points $c_{1},\dots,c_{k}$ and $\ell$, which means $k$ axis-aligned squares centered at $c_{1},\dots,c_{k}$ with side length $\ell$ can cover the polygon
Return : $k$ axis aligned squares, with center points $c_{1},\dots,c_{k}$ and side length $\ell$.

### v2

# On The Computational Complexity of Minimum Aerial Photographs for Planar Region Coverage

@@ Abstract
With the popularity of drone technologies, aerial photography has become prevalent in many daily scenarios such as environment monitoring, structure inspection, law enforcement etc. A central challenge in this domain is the efficient coverage of a target area with photographs that can entirely capture the region, while respecting constraints such as the image resolution, and limited number of pictures that can be taken. This work investigates the computational complexity of covering a simple planar polygon using squares and circles. Specifically, it shows inapproximability gaps of $1.165$ (for squares) and $1.25$ (for restricted square centers) and develops a $2.828$-optimal approximation algorithm, demonstrating that these problems are computationally intractable to approximate. The intuitions of this work can extend beyond aerial photography to broader applications such as pesticide spraying and strategic sensor placement.
@@ Introduction
Consider the scenario when a drone is tasked to take pictures to capture the boundary of some cropland with a gimbal camera attached to it. We can change the zoom factor $\lambda$ of the camera to make the image footprint larger or smaller with the same resolution, e.g. $3000\times 4000$ in 4K resolution. The gimbal camera usually faces downwards when the drone flies high in the sky to take pictures for terrain inspection for applications like agricultural analysis or mapping. With a greater zoom factor of the gimbal camera, the projected footprint on the ground becomes smaller, but the picture taken contains more detail for each part on the ground.
Unfortunately, the computing or disk resource on the drone is usually limited, so let’s assume the drone can only take at most $k$ pictures in one flight, where $k$ is a parameter induced by the drone hardware. We want to ask the question that what is the largest zoom factor $\lambda$ we can give the gimbal camera such that we can still capture all parts of the region with $k$ pictures. Figure  1 shows a coverage result for covering the boundary or interior of a region with circles or squares, where circles can be imagined as footprints of visible regions using downwards facing fisheye cameras or radars, and squares as the specialization of rectangular images taken by regular cameras.
@@
The theoretical aspects of these aerial photography problems strongly relate to the area of computational geometry, where coverage of point sets (aka. geometric pointclouds) is analyzed to a great extent. For example, the $k$-center problem asks that given a set of points, and a number $k$, what is the minimum radius of $k$ circles centered at the subset of these points to cover all the points (toth2017handbook). Similar problems include the $k$-median problem that asks to minimize the summation of the distance from each point to the closest chosen points among the $k$ chosen points, and the $k$-means problem that tries to solve the sum of squared distance from each point to the closest chosen point. These problems, especially $k$-means, can go to high dimensions in machine learning. However, for the real-world coverage problems for aerial photography, the point sets are associated with specific physical meanings, and are usually inside the $3$-dimensional Euclidean space in the simple case, or in $5$-dimension when normal directions are included.
Notably, the majority of the interesting regions to cover in aerial coverage problems have some extent of continuity, for example, a complete structure, a continuous boundary or fence. Geometric shapes like 2D polygons or 3D meshes create more challenges for the usually studied point set coverage problems while making the problems more realistic. In robotics, the continuity characteristic is usually implied, as a robot handling the coverage or inspection tasks moves in a continuous manner, and it makes sense to assume the region to cover is continuous instead of a scattered point set.
Early studies on covering and packing (Hachbaum1985Approximation) provide strong hardness results and approximation scheme for covering points. Later work (feder1988Optimal) further showed $1.88$-inapproximability gap for point sets coverage using $L_{2}$ metric, and $2$-inapprox. gap for coverage using $L_{1}$ metric. For continuous regions and boundaries of a simple polygon, recent work (Feng2020Optimally) has shown $1.152$-inapprox. gap with the minimum circle radius using $k$ circles.
@@
In this paper, the complexity analysis for covering a simple planar polygon using squares is the main focus. Specifically, this work shows the inapproximability gap of $1.165$ for finding the minimum side length of $k$ squares to cover the simple 2D region (a simple polygon) and the inapproximability gap of $1.25$ when the square centers are restricted to be inside the region. This work also develops a $2.828$-optimal algorithm to solve the square coverage problem in polynomial time. Since most regular cameras used by gimbals take rectangular images, the results in this work bring the computational complexity results shown in (Feng2020Optimally) closer to reality for the aerial photography problems. And it can potentially be a reference for setting camera zooming factors or designing gimbal cameras for aerial surveying or mapping.
Additionally, as the drone’s position may be restricted to be only on top of the region to take pictures for the region itself, the paper also considers the variant of the problem when the positions of the square or circle centers must be inside the simply connected region or on top the region’s perimeter. This variant gives the work in (Feng2019Efficient) an aerial photography interpretation.
@@
Given a region, represented by a simple polygon (a planar polygonal chain without self-intersections or holes) $P$, and a number $k\geq 1$. What is the minimum length $\ell$ such that we can put $k$ circles each with radius of $\ell$, such that $P$ can be contained by these circles.
@@
Given a region (a simple polygon $P$), and a number $k\geq 1$. What is the minimum length $\ell$ such that we can put $k$ squares each with side length of $\ell$, such that $P$ can be contained by these squares.
@@
If we only care about covering the boundary of a simple polygon, similar to the definition of the barrier coverage problem (Gage1992Command). This problem can be reduced to the two problems aforementioned because we can solidify the boundary by expanding the edges of a polygon to sticks with a small width $\delta$, and leave a narrow opening on the polygon. When $\delta\rightarrow 0$, the results of coverage problems the solidified boundary skeleton and boundary skeleton converge to be the same. In this way, covering the expanded polygon from the original perimeter is the same as covering the boundary of the original polygon. Conversely, a simple polygon can be approximately represented by a polygon whose boundary almost fills in the interior of the polygon. Hence, the computational complexity of covering the interior and the boundary of a simple region is at the same level.
In many real scenarios, the sensors’ or the robots’ locations (i.e. the centers of the circles or squares) must be inside the region itself or on the perimeter, for example, the watch towers should be built on the defensive wall itself and not outside; unmanned aerial vehicles for securities should not go outside or inside the security zone lines. When the constraint that the circle or square centers must be inside the polygon needs to be satisfied, the constrained version of Problem 2 and  1 is defined as:
Problem 3 (Constrained Circle or Square Coverage)
Same as Problem  2 and Problem  1, but the circle or square centers must stay inside the polygon $P$.
@@
Similarly, the case when the feasible region to place circle centers is the interior of the polygon is equivalent to the case when the feasible region is on the boundary of a polygon, because we can easily make a polygon out of the boundary of a simple polygon by solidifying the line segments of the polygon and making narrow openings for holes elimination.
@@
Given a planar cubic graph $G(V,E)$, we convert it into a special structure through a sequence of steps. The first step embeds the graph $G$ into the plane. Each vertex in $G$ becomes a vertex junction, and each edge becomes a “fence” like structure connecting two neighboring junctions which is an odd length path intersected with a set of perpendicular bars with length $\zeta$ ($\zeta$ is a parameter to be determined for proving different results), at each unit line segment.
@@
From here, the second modification is made which completes the construction of $T_{G}$. In each previously constructed path $u\cdots w=uv_{1}\ldots v_{2m}w$, for each $v_{i}v_{i+1}$, $1\leq i\leq 2m-1$, we add a line segment of length $\zeta$ that is perpendicular to $v_{i}v_{i+1}$ such that $v_{i}v_{i+1}$ and the line segment divide each other in the middle. $G^{\prime}$ and the bars form the intermediate structure $T_{G}$.
@@
The theorem in this section is proved completely in (Feng2020Optimally), but to ease the the following derivations, we give a simplified version of the proof.
@@
The gadget structure (refer it as $T$) constructed in the previous section can be easily converted to a simple polygon by dilating the lines to a thin structure with a small width $\varepsilon$ while leaving a narrow opening with width $\varepsilon$ to make it a simple polygon (refer it as $P_{\varepsilon}$). Assume the minimum radius of $k$ circles to cover a simple polygon is $\ell^{*}$, then the minimum radius $\ell$ for covering $P_{\varepsilon}$ must satisfy $\ell-\varepsilon\leq\ell^{*}\leq\ell+\varepsilon$. When $\varepsilon\rightarrow 0$, $\ell$ and $\ell^{*}$ will converge to be the same. So, we only need to show the complexity of approximation to put $k$ circles with min radius to cover the structure $T$.
@@
Since each junction maps to a vertex in $G$ in the transformation, call the side of the edge with a circle covering only one bar the “odd-end”, and the other side of the edge the “even-end”. If there is a vertex cover of graph $G$, then it is possible to let all the vertices been chosen be the “even-end” or “odd-end”, and the rest of the vertices be “even-end” in the edge link coverage pattern. Since every junction represented vertex is either inside a vertex cover or has a neighbor inside the vertex cover, it is always possible to find such an “odd-end” and “even-end” combination for the junctions so at least no junction has all three edges as even ends.
On the other side, for a coverage of $T$ with the minimum number of circles, select those vertices with at least one even-end. It is clear that these vertices form a vertex cover.
@@
Listing all necessary geometric constraints in Figure  4 gives
@@
Integrating the system of equations from (1) to (6) together gives $\ell=1.152$, which is the extreme case when the circle coverage pattern of $T_{G}$ cannot map to a vertex cover of the original vertex cover problem.
@@
In the structure gadget constructed in the previous section, let us shrink the unit length segment to $\sqrt{2}/2$, so the current distance between neighboring bars becomes ${\sqrt{2}}/{2}$. And the bar length $\zeta$ here is set to be $\zeta={\sqrt{2}}/{2}$.
@@
Figure 5: Structure within the odd length path, where each unit length segment is now shrunk to $\sqrt{2}/2$, and the vertical bar length is $\zeta=\sqrt{2}/2$.
@@
Figure 6: Possible square coverage pattern along the constructed odd-length path until the rectangle side length $\ell$ reaches $1.25$.
@@
Then, let’s prove when there is $n+M$ squares with edge length less than $1.165$ to cover the tri-net structure, there is a vertex cover of size $n$ for the graph $G$.
@@
Figure 8: The pattern of covering the crossing in the special structure constructed at the extreme square side length $\ell$.
@@
At the maximum square side length of $\ell$ such that the structure still needs the same number of squares to cover as using unit squares, Figure  8 shows this situation when the junction can be covered with three squares, among which two squares cover one vertical bar each and one square covers two vertical bars.
Listing all necessary geometric constraints in Figure  8 gives
@@
This inapproximability factor can be interpreted as the maximum zoom factor the gimbal camera can have before taking pictures, or relates to the minimum Ground Sampling Distance (GSD) that $k$ pictures can provide subject to full coverage of the boundary of a simple region using a polynomial time algorithm unless P=NP.
@@
In the construction and also in the problem setting in Problem  1 and  2, it is assumed that the center of the circle or square can be anywhere on the plane. However, in many robotics scenarios, the center of the circles or squares, i.e. the robot locations, has restrictions like they must be on top of the perimeter or inside the region. If those restrictions are applied, are these problems still NP-hard or they can be proved to have a higher inapproximability ratio?
Since if there is a solution of $k$ unit circle to cover the simple polygon constructed, they can be made to be centered at the skeleton structure $T_{G}$ constructed before until the circle radius of $\ell$ reaches $1.152$, the restricted version of the problem has at least the same inapproximability ratio.
@@
At the maximum square side length of $\ell$ such that the structure still needs the same number of squares to cover as using unit squares, Figure  9 shows this situation when the junction can be covered with three squares, among which two squares cover one vertical bar each and one square covers two vertical bars.
Listing all necessary geometric constraints in Figure  9 gives
@@
Integrating the system of equations in (14) to (17) together gives $\ell=1.289$. Considering the possible change of coverage pattern along the odd-length-path depicted in Figure  6 when $\ell$ reaches $1.25$. We have,
@@
Compared to the unrestricted version of the square coverage problem which has the inapproximability ratio of $1.165$, the inapproximability ratio for the restricted version for the square coverage problem is higher by around $8.5\%$.
@@
In a previous work (Feng2020Optimally), sampling followed by applying the traditional $2$-optimal farthest clustering algorithm (Gonzalez1985Clustering) for the $k$-center problem gives a $(2+\epsilon)$-optimal result. When replacing the farthest clustering algorithm with mathematical programming, with the state-of-the-art integer programming tool (Optimization2019Gurobi), a $(1+\epsilon)$-optimal result for $k\leq 100$ can be obtained in $1$ minute on a desktop platform with intel i7 and 16GB memory.
For the square coverage problem, which is similar to the $k$-center problem with $L_{\infty}$ distance metric where the distance between two points $p_{1}=(x_{1},y_{1})$ and $p_{2}=(x_{2},y_{2})$ is $\max(\|x_{1}-x_{2}\|,\|y_{1}-y_{2}\|)$. If we are using $L_{\infty}$ distance metric, which means the square used for covering the target region is always axis-aligned, using the same farthest clustering algorithm (Gonzalez1985Clustering) can produce a 2-optimal solution. Besides, a square with side length $\ell$ facing at any direction can be fully covered by an axis-aligned square with side length $\sqrt{2}\ell$. So, squares created with regular farthest clustering algorithm for the sampled points using the $L_{\infty}$ metric gives a $(2\sqrt{2}+\epsilon)\approx(2.828+\epsilon)$-optimal algorithm.
In the constrained version of the square and circle coverage problems, since the farthest clustering algorithm only selects points (circle or square centers) among the sampled points in the target region or boundary, the results of the previous algorithm automatically satisfy the constraints. Thus, the efficient algorithms for the square and circle coverage problems still apply for approximately solving the constrained circle and square coverage problem within an approximation factor of $2$ and $2.828$, respectively.
@@
1 Sample NN points from the polygon with sampling density $\epsilon$
2 Run the 22-optimal farthest clustering algorithm on the sampled points with $L_{\infty}$ metric to get center points $c_{1},\dots,c_{k}$ and radius $\ell^{\prime}$, which means $k$ axis-aligned squares centered at $c_{1},\dots,c_{k}$ with side length $\ell=2\ell^{\prime}$ can cover the polygon
Return : $k$ axis aligned squares, with center points $c_{1},\dots,c_{k}$ and side length $\ell=2\ell^{\prime}$.